solve the equation by factoring
8x^2 +10x-7 = 0

Answer by LinnW(1048)

8x^2 +10x-7 = 0
product of outer coefficients = -56
we want two numbers whose sum is 10 and product is -56
(-4)(14) = -56
8x^2 +14x -4x -7 = 0
2x(4x + 7) -1(4x + 7) = 0
(2x - 1)(4x + 7) = 0
2x - 1 = 0 ; 2x = 1 ; x = 1/2
4x + 7 = 0 ; 4x = -7 ; x = -7/4